### v1

@@ 3 $k$-Homogeneous ETFs
For any $U$ in the line automorphism group, there exists $\sigma\in S_{n}$ and $(\eta_{j})^{n}_{j=1}\in\operatorname{U}(1)^{n}$ such that
$U\varphi_{j}=\eta_{j}\varphi_{\sigma(j)},\quad j=1,2,\ldots,n.$
By [LW65, Theorem 2A], if $n\geq 2k$, then any $k$-homogeneous group action on a set of size $n$ is $(k-1)$-transitive. Thus, if $n\geq 4$, then the doubly homogeneous line automorphism group must act $1$-transitively, meaning that the vectors are all equal norm. If $n=3$, then the doubly homogeneous line automorphism group must act $(3-2)$-homogeneously, i.e., $1$-transitively, by Lemma 3.3, and again the vectors must be equal norm.
Let $j\neq k$ and $\ell\neq m$. Since the line automorphism group is doubly homogeneous, there exists a $U\in\operatorname{U}(d)$ with an associated permutation $\sigma$ yielding $\{\sigma(j),\sigma(k)\}=\{\ell,m\}$ as sets. Since $j,k,\ell,m$ were arbitrary, we have that there exist $0\leq\alpha\leq\beta$ such that for any $j,k$, $\left|\langle\varphi_{j},\varphi_{k}\rangle\right|=(\beta-\alpha)\delta_{j,k}+\alpha$.
A frame is tight if and only if $\Phi\Phi^{\ast}=AI$ if and only if the Gram matrix $\Gamma=\Phi^{\ast}\Phi$ is $A$ times an orthogonal projection. This is the formulation we will use. Let $j,k,\ell,m$ be as above. Then
$\Gamma_{j,k}$ $=\left\langle\varphi_{k},\varphi_{j}\right\rangle=\left\langle U\varphi_{k},U\varphi_{j}\right\rangle$
$=\left\langle\eta_{k}\varphi_{\sigma(k)},\eta_{j}\varphi_{\sigma(j)}\right\rangle$
$=\eta_{k}\overline{\eta_{j}}\left\langle\varphi_{\sigma(k)},\varphi_{\sigma(j)}\right\rangle$
$=\overline{\eta_{j}}\eta_{k}\Gamma_{\sigma(j),\sigma(k)}$
=$=\left\{\begin{array}[]{c}\overline{\eta_{j}}\eta_{k}\Gamma_{m,\ell}\\[2.0pt] \overline{\eta_{j}}\eta_{k}\Gamma_{\ell,m}\end{array}\right.$
Similarly,
$\left(\Gamma^{2}\right)_{j,k}$ $=\sum_{i=1}^{n}\left\langle\varphi_{i},\varphi_{j}\right\rangle\left\langle\varphi_{k},\varphi_{i}\right\rangle$
$=\sum_{i=1}^{n}\left\langle U\varphi_{i},U\varphi_{j}\right\rangle\left\langle U\varphi_{k},U\varphi_{i}\right\rangle$
$=\sum_{i=1}^{n}\left\langle\eta_{i}\varphi_{\sigma(i)},\eta_{j}\varphi_{\sigma(j)}\right\rangle\left\langle\eta_{k}\varphi_{\sigma(k)},\eta_{i}\varphi_{\sigma(i)}\right\rangle$
$=\overline{\eta_{j}}\eta_{k}\sum_{i=1}^{n}\left\langle\varphi_{\sigma(i)},\varphi_{\sigma(j)}\right\rangle\left\langle\varphi_{\sigma(k)},\varphi_{\sigma(i)}\right\rangle$
$=\overline{\eta_{j}}\eta_{k}\left(\Gamma^{2}\right)_{\sigma(j),\sigma(k)}$
=$=\left\{\begin{array}[]{c}\overline{\eta_{j}}\eta_{k}\left(\Gamma^{2}\right)_{m,\ell}\\[2.0pt] \overline{\eta_{j}}\eta_{k}\left(\Gamma^{2}\right)_{\ell,m}\end{array}\right.$
Thus, the off-diagonal of $\Gamma^{2}$ is a constant multiple of the off-diagonal of $\Gamma$. Also,
$\left(\Gamma^{2}\right)_{j,j}=\sum_{i=1}^{n}\left|\langle\varphi_{i},\varphi_{j}\rangle\right|^{2}=\beta^{2}+(n-1)\alpha^{2}=\frac{\beta^{2}+(n-1)\alpha^{2}}{\beta}\Gamma_{j,j}$
Hence, there exists $c_{1},c_{2}\in\mathbb{C}$ such that $\Gamma^{2}=c_{1}\Gamma+c_{2}I$. Since $n>d$, the Gram matrix $\Gamma$ cannot be invertible, meaning that $c_{2}=0$. (Otherwise $(\Gamma-c_{1}I)/c_{2}$ would be the inverse.) But this means that $\Gamma^{2}=c_{1}\Gamma$, meaning $\Gamma$ is a scalar multiple of an orthogonal projection and hence $\Phi$ is a tight frame. Putting everything together, we have that $\Phi$ is an ETF. ∎
In a sense, $2$-homogeneous ETFs are the only interesting $k$-homogeneous ETFs to study. $k$-homogeneous ETFs for $2<k\leq 5$ are either $2$-transitive (and thus one could use those tools to study them, like [IM22, IM24]) or in some sense either trivial or rare. We make this more explicit in what follows. The $(2,4)$-SIC is the (unique up to equivalence) ETF of $4$ vectors in $\mathbb{C}^{2}$.
Theorem 3.6.
The only ETFs with $3$-homogenous line automorphism groups are the $(2,4)$-SIC, the $(d,2d)$-ETFs with $d=q+1$ where $q\equiv 3\bmod 4$ is a prime power arising from Paley skew-conference matrices, orthonormal bases, and simplices, which are all $2$-transitive. The only ETFs with $4$-homogenous line automorphism groups are the $(2,4)$-SIC, the $(3,6)$ real ETF, orthonormal bases, and simplices. The only ETFs with $5$-homogenous line automorphism groups are the $(3,6)$ real ETF, the $(3,7)$- and $(4,7)$-ETFs (at least of Paley type), the $(4,8)$-ETF, orthonormal bases, and simplices.
@@
We consider $(d,n)$-ETFs, using [FM16] as a reference for which $(d,n)$ yield ETFs. By [LW65], if $n\geq 2k$, then any $k$-homogeneous group action on a set of size $n$ is $(k-1)$-transitive. Thus, all $3$-homogeneous ETFs (lines or vectors) with $n\geq 6$ are also $2$-transitive. By Lemma 3.4, the only ETFs which are not simplices or orthonormal bases but have $3$-homogeneous line automorphism group must have $n=2d$. It follows from [IM24] that the only $2$-transitive ETFs with $n=2d$ and purely imaginary inner products have $d=q+1$ with $q\equiv 3\bmod 4$ and are generated by Paley skew-conference matrices. The only ETF with $n<6$ which is not an orthonormal basis or simplex is the $(2,4)$ SIC, which is $1$-homogeneous as a SIC and thus $3=(4-1)$-homogeneous as well. Note that the $(2,4)$-SIC has $2$-transitive line automorphism group [Zhu15]; thus, all of the ETFs with $3$-homogeneous line automorphism group are also $2$-transitive.
Applying [LW65] again, we see that all $4$-homogeneous ETFs (lines or vectors) with $n\geq 8$ are also $3$-transitive, which means they are simplex ETFs or orthonormal bases (Lemma 3.2). Similarly, ETFs with $84\leq n<$ 8 which are orthonormal bases or simplex ETFs trivially have $4$-homogeneous line automorphism groups (Lemma 3.1). The remaining ETFs which have $n<8$ are the $(2,4)$-SICs, the $(3,6)$ real ETF (which is unique up to equivalence in $\mathbb{R}^{3}$ [BK06]), the $(3,7)$-ETF(s) and $(4,7)$-ETF(s). The $(2,4)$ SIC is trivially $4$-homogeneous since the identity maps the group of four lines to the group of four. The $(3,6)$ real ETF corresponds to the antipodal vertices of the regular icosahedron and thus has line automorphism group isomorphic to $\operatorname{PSL}(2,5)$, which acts $2$ transitively on the lines (see, e.g., Theorem 1.3 [IM24] which also gives uniqueness of the lines), in turn implying that the automorphism group is $2$-homogeneous and thus $4=(6-2)$-homogeneous. The remaining ETF parameters are the $(3,7)$- and $(4,7)$-ETFs. We have already shown that neither is $3$-homogeneous; thus, neither can have a $4=(7-3)$-homogeneous line automorphism group.
We now analyze ETFs for $5$-homogeneous line automorphism groups. As before, we only need to consider ETFs with $n<10=2(5)$. We have already shown that $(3,9)$ and $(6,9)$-ETFs cannot be $4$-homogeneous; thus, they are not $5=(9-4)$-homogeneous. Next we consider the $(4,8)$ ETF (which is $2$-transitive and unique up to equivalence [IM24]). In the standard construction, the inner products are all purely imaginary; thus, Lemma 3.4 implies that the ETF is $3$-homogeneous and in turn $5=(8-3)$-homogeneous. The $(3,7)$-Paley ETF and its Naimark complement are $2$-homogeneous, hence $5=(7-2)$-homogeneous. Finally, the $(3,6)$ real ETF is $1$-transitive, hence $1$-homogeneous and $5=(6-1)$-homogeneous. ∎
In the remaining cases where $k>5$, any $k$-homogeneous ETF with $n\geq 2k$ will be $(k-1)$-transitive and thus can only be a simplex ETF or orthonormal basis. For $n<2k$, the ETF will be $(n-k)$-homogeneous with $n-k<k$. Thus, inductively, the only possibilities are $1$-homogenous ETFs (i.e., the ETF lines are spun up as the orbit of a group with no additional symmetries), $2$-homogeneous ETFs, or any of the $3$, $4$, or $5$-homogeneous ETFs enumerated above.
Corollary 3.7.
The only ETFs with $3$-homogenous vector automorphism groups are orthonormal bases and the regular simplex ETFs. The only ETFs with $4$-homogenous vector automorphism groups are the $(2,4)$-SIC, orthonormal bases, and simplices.
Proof.
By following the cases in the proof of Theorem 3.6, we see that the only possible ETF which has a $3$-homogeneous vector automorphism group which is not doubly transitive is the $(2,4)$-SIC. Although any SIC is the orbit of a fiducial under a unitary system, the unitary system is a projective unitary representation and thus does not act transitively on the vectors. Hence, the vector automorphism group of the $(2,4)$-SIC is neither $1$-homogeneous nor $3=4-1$-homogeneous (Lemma 3.3). All other ETFs with ETFs with $3$-homogenous line automorphism groups must be doubly transitive, and hence Lemma 3.2 applies.
By Theorem 3.6, we just need to check the $(2,4)$-SIC and the $(3,6)$ real ETF for $4$-homogeneous automorphism vector groups. The $(2,4)$ SIC is trivially $4$-homogeneous. If the $(3,6)$ real ETF had a $4$-homogeneous vector automorphism group, then (Lemma 3.3) it would be $2=(6-4)$-homogeneous, which (Lemma 3.2 cannot happen since it is neither an orthonormal basis or a regular simple ETF. ∎
@@ 4 Dave Larson’s AMS Memoirs
Assume that $\Phi=(\varphi_{j})_{j\in J}$ is a tight frame with frame bound $A$ for an infinite dimensional separable Hilbert space $\mathcal{H}$. Then $J$ is countably infinite. If $\Phi$ were doubly homogeneous, then $\left|\langle\varphi_{i},\varphi_{j}\rangle\right|=\left|\langle\varphi_{k},\varphi_{\ell}\rangle\right|$ for all $i\neq j$ and $k\neq\ell$. Call this number $\alpha$. By the definition of a tight frame (Definition 1.1), we see that
$\left\|\varphi_{k}\right\|^{2}=\frac{1}{A}\sum_{j\in J}\left|\langle\varphi_{k},\varphi_{j}\rangle\right|^{2}=\frac{1}{A}\left(\left\|\varphi_{k}\right\|^{2}+\sum_{j\neq k}\alpha^{2}\right).$
Since the left-hand-side is a finite number, the right-hand-side must converge, meaning that $\alpha=0$. Thus, the only infinite doubly homogeneous infinite frames are orthogonal bases.
@@ 4.1 Problems
When $q\equiv 3\bmod 4$ is a prime power $q=p^{s}$, one may still create a Paley ETF $\Phi_{q}$ by pulling out rows corresponding to the quadratic residues. As in Example 1.12, the frame vectors may be expressed as an orbit of $\mathbbm{1}$, i.e., $\varphi_{j}=M_{j}\mathbbm{1}$ for $M_{j}$ the $d\times d$ diagonal matrix with diagonal entries the entries of $\varphi_{j}$. However, when $s>1$, $\{M_{j}\}_{j}$ is a non-cyclic group (a representation of the non-cyclic additive group of $\mathbb{F}_{q}$). A certain group of translations will also be in the line and vector automorphism groups. In code from Joey Iverson, Harley Meade computed that the vector automorphism group of $\Phi_{3^{3}}$ is $(\mathbb{Z}_{3})^{3}\rtimes\mathbb{Z}_{3}$ while the line automorphism group is $\left((\mathbb{Z}_{3})^{3}\rtimes\mathbb{Z}_{13}\right)\rtimes\mathbb{Z}_{3}$. Calculations for larger non-cyclic Paley ETFs proved to be too unwieldy.
@@
What are the vector and line automorphism groups of the Paley ETFs $\Phi_{q}$ with non-cyclic additive groups?
As referenced throughout this paper, various combinatorial designs arising from Paley-like constructions have reasonably high amounts of symmetry. It seems reasonable to assume that the non-cyclic Paley ETFs will be doubly homogeneous. We know that they are not doubly transitive [IM22, IM24]. By [Kan69], any set of size $n\geq 4$ which has a doubly homogeneous but not doubly transitive automorphism group must have $n=q$ for $q\equiv 3\bmod 4$ a prime power $q=p^{s}$ with the automorphism group isomorphic to a subgroup of
$\operatorname{A\Gamma L}(\mathbb{F}_{q})=\left\{\,f:\mathbb{F}_{q}\rightarrow\mathbb{F}_{q}\,\,\middle|\,\,f(x)=mx^{\sigma}+b,\,\,m\in\mathbb{F}_{q}^{\times},\,\,\sigma\in\operatorname{Aut}(\mathbb{F}_{q}),\,\,b\in\mathbb{F}_{q}\,\right\}$
and isomorphically containing
$\left\{\,f\in\operatorname{A\Gamma L}(\mathbb{F}_{q})\,\,\middle|\,\,f(x)=x+b,\,\,b\in\mathbb{F}_{q}\,\right\}.$
Putting all of this together, we conjecture an answer to Question 4.1.
Conjecture 4.2.
Let $\Phi_{q}$ be a Paley ETF with $q=p^{s}$. Then $\Phi_{q}$ has a line automorphism group isomorphic to
$\left(\left(\mathbb{Z}_{p}\right)^{s}\rtimes\mathbb{Z}_{(q-1)/2}\right)\rtimes\mathbb{Z}_{s}\cong\left\{\,f\in\operatorname{A\Gamma L}(\mathbb{F}_{p})\,\,\middle|\,\,f(x)=mx^{\sigma}+b,\,\,m\in\operatorname{QR}(\mathbb{F}_{q}),\,\,\sigma\in\operatorname{Aut}(\mathbb{F}_{q}),\,\,b\in\mathbb{F}_{q}\,\right\}$
Due to a result from Chebotarëv [SL96], the Paley ETFs $\Phi_{p}$ with $p$ prime are full spark, i.e., have spark $(p-1)/2+1=(p+1)/2$. Their bender is just trivially the set of all subsets of size $(p+1)/2$, which is a $t$-design for all $21\leq t\leq(p+1)/$ 2. Calculations from Harley Meade show that the spark of $\Phi_{3^{3}}$ is $8$ and the bender is a $(27,8,28)$-BIBD. Again, calculations for larger non-cyclic Paley ETFs were not tractable.
Question 4.3.
@@
We note that the computation of the Gram matrix of $\Phi_{p}$ for $p$ prime in the proof of Lemma 2.6 shows that the $\Phi_{p}$ are neither of two other categories of ETFs which are in some sense doubly-transitive-adjacent, i.e., being DRACKN or roux [CGSZ16, IM22]. This leads to the following question.
Question 4.5.
Are there any ETFs which are simultaneously doubly homogeneous and DRACKN or doubly homogeneous and roux but not doubly transitive?
@@
Question 4.6.
Are the elements of the automorphism group of a Fourier ETF (or, more generally, a Fourier frame) always generalized permutations?
@@ Acknowledgments
The author is especially grateful to Harley Meade, who ran code to help formulate hypotheses and explicitly proved that a group related to $\{M^{k}T^{\ell}\}$ acts doubly homogeneously but not doubly transitively on the Naimark complement of a prime order Paley ETF (which was replaced with Proposition 3.1 in [Kan69] in the proof of Theorem 2.8). The author would also like to thank Joey Iverson and Dustin Mixon for fruitful discussions and also thank Joey for providing GAP code which allowed the exploration of various examples. Most of all, the author is grateful to Dave Larson for introducing her to frames twenty-something years ago.

### v2

@@ 3 $k$-Homogeneous ETFs
We consider $(d,n)$-ETFs, using [FM16] as a reference for $(d,n)$ which yield ETFs. By [LW65], if $n\geq 2k$, then any $k$-homogeneous group action on a set of size $n$ is (k1)(k-1)-transitive. Thus, all $3$-homogeneous ETFs (lines or vectors) with $n\geq 6$ are also $2$-transitive. By Lemma 3.4, the only ETFs which are not simplices or orthonormal bases but have $3$-homogeneous line symmetry group must have $n=2d$. It follows from [IM24] that the only $2$-transitive ETFs with $n=2d\geq 6$ and purely imaginary inner products have $d=q+1$ with $q\equiv 3\bmod 4$ and are generated by Paley skew-conference matrices. These have line symmetry group $\operatorname{PSL}(2,q)$ acting necessarily faithfully and transitively, where (see, e.g., [Kan72]) $\operatorname{PSL}(2,q)$ acts $3$-homogeneously. The only ETF with $n<6$ which is not an orthonormal basis or simplex is the $(2,4)$-SIC, which is $1$-homogeneous as a SIC and thus 3=(41)3=(4-1)-homogeneous as well. Note that the $(2,4)$-SIC has $2$-transitive line symmetry group [Zhu15]; thus, all of the ETFs with $3$-homogeneous line symmetry group are also $2$-transitive.
Applying [LW65] again, we see that all $4$-homogeneous ETFs (lines or vectors) with $n\geq 8$ are also $3$-transitive, which means they are simplex ETFs or orthonormal bases (Lemma 3.2). Similarly, ETFs with $4\leq n<8$ which are orthonormal bases or simplex ETFs trivially have $4$-homogeneous line symmetry groups (Lemma 3.1). The remaining ETFs which have $n<8$ are the $(2,4)$-SIC, the $(3,6)$-ETFs, the $(3,7)$-ETF(s) and $(4,7)$-ETF(s). The $(2,4)$ SIC is trivially $4$-homogeneous since the identity maps the group of four lines to the group of four. If a (not necessarily real) $(3,6)$-ETF is $4$-homogeneous, then it is $2$-homogeneous. Since $6$ is not prime power, any $2$-homogeneous $(3,6)$-ETF must be $2$-transitive [Kan69b]. The real $(3,6)$-ETF corresponds to the antipodal vertices of the regular icosahedron and thus has line symmetry group isomorphic to $\operatorname{PSL}(2,5)$, which acts $2$ transitively on the lines (see, e.g., Theorem 1.3 [IM24] which also gives uniqueness up to equivalence of the $(3,6)$-ETFs with $2$-transitive symmetry), in turn implying that the automorphism group is $2$-homogeneous and thus 4=(62)4=(6-2)-homogeneous. The remaining ETF parameters are the $(3,7)$- and $(4,7)$-ETFs. We have already shown that neither is $3$-homogeneous; thus, neither can have a 4=(73)4=(7-3)-homogeneous line symmetry group. ∎
As with the characterization of $4$-homogeneity, in the remaining cases where $k>4$, any $k$-homogeneous ETF with $n\geq 2k$ will be (k1)(k-1)-transitive and thus can only be a simplex ETF or orthonormal basis. For $n<2k$, the ETF will be (nk)(n-k)-homogeneous with nk<kn-k<k. Thus, inductively, the only possibilities are $1$-homogeneous ETFs (i.e., the ETF lines are spun up as the orbit of a group with no additional symmetries), $2$-homogeneous ETFs, or any of the $3$ or $4$-homogeneous ETFs enumerated above.
There are skew-conference matrices which are not Paley (e.g., [CD07]). Thus, being an ETF with purely imaginary triple products is not sufficient to be $3$-homogeneous.
Corollary 3.6.
The only ETFs with $3$-homogeneous vector symmetry groups are orthonormal bases and the regular simplex ETFs. The only ETFs with $4$-homogeneous vector symmetry groups are the $(2,4)$-SIC, orthonormal bases, and simplices.
@@
By following the cases in the proof of Theorem 3.5, we see that the only possible ETF which has a $3$-homogeneous vector symmetry group which is not $2$-transitive is the $(2,4)$-SIC. Although any SIC is the orbit of a fiducial under a unitary system, the unitary system is a projective unitary representation and thus does not act transitively on the vectors. Hence, the vector symmetry group of the $(2,4)$-SIC is neither $1$-homogeneous nor 3=(41)3=(4-1)-homogeneous (Lemma 3.3). All other ETFs with ETFs with $3$-homogeneous line symmetry groups must be $2$-transitive, and hence Lemma 3.2 applies.
By Theorem 3.5, we just need to check the $(2,4)$-SIC and the $(3,6)$ real ETF for $4$-homogeneous automorphism vector groups. The $(2,4)$-SIC is trivially $4$-homogeneous. If the $(3,6)$ real ETF had a $4$-homogeneous vector symmetry group, then (Lemma 3.3) it would be 2=(64)2=(6-4)-homogeneous, which (Lemma 3.2) cannot happen since it is neither an orthonormal basis or a regular simple ETF. ∎
@@ 4 Dave Larson’s AMS Memoirs
The preceding results mirror several themes in [HL00, DL98]; namely, that group actions are a powerful tool to construct frames with desirable properties, that Naimark complementation can be used to simplify the search for frames with particular properties, and that associated groups other than those used to generate the frame are powerful tools in understanding the frame properties.
@@
Assume that $\Phi=(\varphi_{j})_{j\in J}$ is a tight frame with frame bound $A$ for an infinite dimensional separable Hilbert space $\mathcal{H}$. Then $J$ is countably infinite. If $\Phi$ were $2$-homogeneous, then $\left|\langle\varphi_{i},\varphi_{j}\rangle\right|=\left|\langle\varphi_{k},\varphi_{\ell}\rangle\right|$ for all $i\neq j$ and $k\neq\ell$. Call this number $a$. By the definition of a tight frame (Definition 1.1), we see that
$\left\|\varphi_{k}\right\|^{2}=\frac{1}{A}\sum_{j\in J}\left|\langle\varphi_{k},\varphi_{j}\rangle\right|^{2}=\frac{1}{A}\left(\left\|\varphi_{k}\right\|^{2}+\sum_{j\neq k}a^{2}\right).$
Since the left-hand-side is a finite number, the right-hand-side must converge, meaning that $a=0$. Thus, the only infinite $2$-homogeneous infinite frames are orthogonal bases.
@@ 4.1 Problems
Corollary 2.20 gives that the bender of any Paley ETF is a BIBD. However, it does not give information on the parameters of the BIBD. The bender of a Paley ETF $\Phi_{p}$ for $p$ a prime is simply the set of all subsets of size $(p+1)/2$. Calculations from Harley Meade show that the spark of $\Phi_{27}$ is $8$ and the bender is a $(27,8,28)$-BIBD. Calculations for Paley ETFs with larger $q$ were not tractable.
@@
To aid in solving the conjecture, there is code [Kin] to allow one to generate Paley ETFs of arbitrary size. This code also generates the symmetries $T$ and $\Pi$. There is additionally a function to construct ETFs from any difference sets.
We note that the computation of the Gram matrix of $\Phi_{p}$ for $p$ prime in the proof of Lemma 2.11 shows that the $\Phi_{p}$ are neither of two other categories of ETFs which are in some sense $2$-transitive-adjacent, i.e., being DRACKN or roux [CGSZ16, IM22]. This leads to the following question.
Question 4.3.
Are there any ETFs which are simultaneously $2$-homogeneous and DRACKN or $2$-homogeneous and roux but not $2$-transitive?
@@
Question 4.4.
Are the elements of the automorphism group of a Fourier ETF always generalized permutations?
We point to code from Joey Iverson [Ive] to determine the vector and line symmetry groups of frames as a resource to help answer this question.
@@ Acknowledgments
The author is especially grateful to Joey Iverson, for fruitful discussions and GAP code [Ive] to run some experiments. The author is also thankful to Harley Meade, who ran Joey’s code to help formulate hypotheses and explicitly proved that groups related to $\{M^{k}T^{\ell}\}$ act $2$-homogeneously on ETFs related to prime-order Paley ETFs, to Dustin Mixon for helpful feedback, and to the anonymous referees for their thoughtful suggestions which lead to great improvements in the text. Most of all, the author is grateful to Dave Larson for introducing her to frames twenty-something years ago.